OpenStudy (anonymous):

Anne and Nancy use a metal alloy that is 15.6% copper to make jewelry. How many ouncesw of an alloy that is 14% copper must be mixed with an alloy that is 16% copper to form 100 ounces of the desired alloy?

OpenStudy (tkhunny):

Equate Stuff. F = Amount of 14% Copper S = Amount of 16% Copper f + S = 100 Equate Copper F*0.14 + S*0.16 = 100*0.156 Equate NonCopper F*0.86 + S*0.84 = 100*0.844

OpenStudy (anonymous):

I'm sorry, I'm not seeing how to solve this. What is the equation I'm supposed to use?

OpenStudy (tkhunny):

You must define variables that will be useful. I did that for you with F and S. You must decide how to solve it - which part to equate. I gave you the two ways for this problem statement. The point of this problem is for yuo to learn how to set it up. The remaining algebra is the easy part.